OpenStudy (anonymous):

Not sure where to start here! Complete the square and find the vertex form of: g(x)= x^2-2x-5

OpenStudy (waleed_imtiaz):

To find the vertex... ? well, to do completing square just add and subtract (1)^2.......

OpenStudy (anonymous):

Ensure that the coefficient of x^2 is 1

OpenStudy (anonymous):

x^2-2x-5 = x^2-2x + (-1)^2 -5 - (-1)^2 = (x - 1)^2 -5 - 1 = (x - 1)^2 - 6

OpenStudy (anonymous):

The vertex is (1, -6)